OpenStudy (anonymous):

What is the solution to the system? Line 1 x y –1 9 3 –3 Line 2 x y –1 –1 2 5 ( _ , _ )

OpenStudy (anonymous):

use the equation of a line formula y=mx+b

OpenStudy (anonymous):

where m is the slope and b is the y-intercept

OpenStudy (anonymous):

find the equation for the first line

OpenStudy (anonymous):

Ok. I'm working on it

OpenStudy (anonymous):

@10115658 Y= -3x + 9

OpenStudy (anonymous):

@10115658 Thank you. I think I have figured it out.

OpenStudy (anonymous):

sorry, i was afk. give me a sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got y=-3x+6

OpenStudy (anonymous):

How did you get the 6?

OpenStudy (anonymous):

y=-3x+b plug in a point and solve for b

OpenStudy (anonymous):

9=-3(-1)+b

OpenStudy (anonymous):

9=3+b b=6

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

I think I got it, b/c I lost you for a sec

OpenStudy (anonymous):

ok, tell me what your answer is and i'll see what i get

OpenStudy (anonymous):

ok, I figured it out

OpenStudy (anonymous):

I'm trying to solve still.

OpenStudy (anonymous):

I think it's 3. What do you think it is?

OpenStudy (anonymous):

what is 3?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

you should get a point x,y

OpenStudy (anonymous):

thats not what i got for x

OpenStudy (anonymous):

I'll try again with a point.

OpenStudy (anonymous):

Do you want me to show you how i got to my answer?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

ok, so i found the equation for the first line as y=-3x+6, which i showed you previously

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

the equation for the second line is y=2x+1, using the same method as the equation for the first line

OpenStudy (anonymous):

now we need to find what x and y are

OpenStudy (anonymous):

so we replace y in the first equation with 2x+1 and solve for x

OpenStudy (anonymous):

2x+1=-3x+6

OpenStudy (anonymous):

Ok I was trying to figure out x and y too.

OpenStudy (anonymous):

solving for x i got x=1

OpenStudy (anonymous):

Then replace x with 1 in one of the equations to find y

OpenStudy (anonymous):

y=2(1)+1

OpenStudy (anonymous):

y=3

OpenStudy (anonymous):

so the point is (1,3)

OpenStudy (anonymous):

Do you agree with the answer?

OpenStudy (anonymous):

Ok, but wouldn't be 9 though,b/c of the y-intercept.

OpenStudy (anonymous):

B I mean

OpenStudy (anonymous):

what do you mean? are you talking about the equation for the first line?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok, theres an easy way to see if the equation is correct or not

OpenStudy (anonymous):

so you think the equation is y=-3x+9 correct?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

ok, so you plug in a point and see if both sides of the equation are equal

OpenStudy (anonymous):

lets plug in -1,9

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

9=-3(-1)+9

OpenStudy (anonymous):

does 3(-1)+9 equal 9?

OpenStudy (anonymous):

Not equal

OpenStudy (anonymous):

so that equation is incorrect lets try my equation

OpenStudy (anonymous):

9=-3(-1)+6

OpenStudy (anonymous):

Yeah lets try it

OpenStudy (anonymous):

9=9 so the equation works

OpenStudy (anonymous):

Yeah the equation works

OpenStudy (anonymous):

so you understand now?

OpenStudy (anonymous):

Yes thank you for helping me. I gave you a medal and became a fan.

OpenStudy (anonymous):

thats a nice visual representation

OpenStudy (anonymous):

It is. It helped.

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

Yeah thanks. I appreciate it.